# custom_references.bib

### v1

@@ 2.5 Constructing Coin Flipping Protocols
We introduce necessary preliminaries in Section 3. In Section 4, we prove our main technical result on jointly biasing most functions in any family of functions. We prove our main lower bound result on the resilience of coin flipping protocols in Section 5. In Section 6, we present new coin flipping protocols. We conclude with some discussion and open problems in Section 7.
@@ 3.1 Notation and Terminology
For $20<\varepsilon<1/$ 2 and $o\in\{0,1\}$, we say that $S\subset[\ell]$ can $(1-\varepsilon)$-bias a function $f:\{0,1\}^{\ell}\to\{0,1\}$ towards $o$ if there exists an adversary $\mathcal{A}_{o}:\{0,1\}^{\ell-\left\lvert S\right\rvert}\to\{0,1\}^{\left\lvert S\right\rvert}$ such that the function $g_{o}(y):=f(\mathcal{A}_{o}(y),y)$ has $\Pr_{\mathbf{y}}[g_{o}(\mathbf{y})=o]\geq 1-\varepsilon$. We will similarly use the notion that a set $S\subseteq[\ell]$ can $(1-\varepsilon)$-bias a protocol $\pi$ towards $o$.
@@
For a function $f:\{0,1\}^{\ell}\to\{0,1\}$, we sometimes write “$\Pr[f=o]$” to denote $\Pr_{\mathbf{y}\sim\{0,1\}^{\ell}}[f(\mathbf{y})=o]$. Similarly, for $\pi$ a $k$-round protocol over $\ell$ players in which each player outputs one bit per round, we sometimes write “$\Pr[\pi=o]$” to denote $\Pr_{\mathbf{x}\sim{(\{0,1\}^{\ell})^{k}}}[\pi(\mathbf{x})=o]$.
@@ 3.2.2 Helpful Claims
3.3 Leader Election and Collective Coin Flipping Protocols
3.3.1 Protocols in the full information model
We formalize the definition of protocols in the full information model. Collective coin flipping protocols and leader election protocols are special cases of such protocols where the output domain is $\{0,1\}$ and $[\ell]$ respectively.
Definition 3.5 (Protocol in the full information model).
A $k$-round protocol with output domain $Y$ over $\ell$ players where each player sends $r$ random bits per round is a function
@@ 3.3.1 Protocols in the full information model
Here is how the protocol operates in the presence of a set $B\subset[\ell]$ of bad players: In round $i$, each of the players from $[\ell]\setminus B$ independently outputs a uniformly random element from $\{0,1\}^{r}$. Let their collective outputs be $\alpha_{i}\in\left(\{0,1\}^{r}\right)^{[\ell]\setminus B}$. Then, depending on $\alpha_{1},\dots,\alpha_{i}$, the players in $B$ together output an element of ($\left(\{0,1\}^{r}\right)^{B}$. Hence, we model the strategy of the bad players as a sequence of functions $\sigma=(\sigma_{1},\dots,\sigma_{k})$, where
@@
In this paper we will mostly consider the case in which $r=1$ and this will be the default assumption unless stated otherwise. Note that when $k=1$, the protocol $\pi$ just becomes a function over $\{0,1\}^{\ell}$; such 1-round coin flipping protocols which cannot be biased by any small set of bad players are also known as resilient functions.
@@
A leader election protocol $\pi$ is a protocol in the full information model with output domain $Y=[\ell]$, the number of players the protocol is operating on. Furthermore, we say $\pi$ is $(b,\gamma)$-resilient if in the presence of any set $B$ of bad players with $\left\lvert B\right\rvert\leq b$, we have that $\Pr[\pi|_{B}\in B]\leq 1-\gamma$.
@@ 3.3.2 A Useful Claim
Fix $\gamma\in(0,1/2]$. Suppose that for every $k$ round coin flipping protocol $\pi^{\prime}$ over $\ell$ players, there exist $b$ bad players that can $(1-\gamma)$-bias $\pi$ towards a particular outcome $o\in\{0,1\}$. Then, for every $(k-1)$-round leader election protocol $\pi$, there exist $b$ bad players such that that a good player is elected as a leader with probability at most $2\gamma$.
@@ 4 Biasing a Family of Functions
In this section, we show that for any family of functions mapping $\{0,1\}^{\ell}\to\{0,1\}$, there exists a “common set” $B_{R}$ of coordinates in $[\ell]$ such that for almost every function $f$ in the family, $B_{R}$ along with a small “heavy set” of coordinates (which may depend on $f$) can together bias $f$. Moreover, crucially, neither $B_{R}$ nor the “heavy set” is too large. This result plays an essential role in our main lower bound for $k$-round coin flipping protocols, just as its simplified version, Theorem 2.2, played a crucial role in the sketch of our two-round lower bound given in Section 2.
@@
Let $G$ be the bipartite graph with left vertex set $U$ given by $\mathcal{F}$ and right vertex set $V$ consisting of all size-$r$ subsets of $[\ell]$, where $r=\frac{100\ell\log(1/\delta)}{\gamma\log(h)}$. $G$ contains an edge between $f\in U$ and $B_{R}\in V$ if there exists $B_{H}=B_{H}(f,B_{R})\subset[\ell]$ with $\left\lvert B_{H}\right\rvert\leq h$ such that $B_{R}\cup B_{H}$ can $(1-\gamma)$-bias $f$ towards $o_{f}$. By Lemma 4.2, the minimum left degree of this graph is at least $(1-\delta)\binom{\ell}{r}$. This implies that the average right degree of this graph is at least $(1-\delta)\left\lvert\mathcal{F}\right\rvert$, and hence, there exists $B_{R}^{*}\in V$ with degree at least $(1-\delta)\left\lvert\mathcal{F}\right\rvert$. Thus, there must exist some fixed $B_{R}^{*}$ such that for $(1-\delta)$ fraction of $f\in\mathcal{F}$, there exists $B_{H}=B_{H}(f)$ with $|B_{H}|\leq h$ such that $B_{R}^{*}\cup B_{H}$ can $(1-\gamma)$-bias $f$ towards $o_{f}$. ∎
@@ 4.1 Proving the Main Lemma
We will use the following concentration inequality regarding submartingales from [RSZ02leader]:
Lemma 4.3 (Lemma 9 of [RSZ02leader]).
@@
We will also use the following result from [RSZ02leader], which follows as a slight extension of [KKL88influence]:
@@
Let $0<\gamma<{\frac{1}{2}},0<\theta<\frac{1}{8},\ell\in\mathbb{N}$. Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be such that $\gamma\leq\Pr[f=1]\leq 1-\gamma$. If $I_{i}(f)\leq\theta$ for all $i\in[\ell]$, then
$\sum_{i=1}^{\ell}I_{i}(f)\geq\frac{\gamma\log(1/\theta)}{20}.$
@@
$\mathbf{X}_{j}=\begin{cases}I_{v_{j}}(f|_{v_{1},\dots,v_{j-1}})&\text{if~{}}\Pr[f|_{v_{1},\dots,v_{j-1}}=1]<1-\gamma\\ 1&\textrm{otherwise}\end{cases}$
@@ 5 Biasing Coin Flipping Protocols
There exist universal constants $C=10^{7},\ell_{0}\in\mathbb{N}$ such that for all $k\in\mathbb{N},\ell\in\mathbb{N},0<\gamma<1/4$, the following holds: For any $k$-round coin flipping protocol $\pi$ over $\ell\geq\ell_{0}$ players where $\Pr[\pi=1]\geq\gamma$, there exists a set of bad players $B\subset[\ell]$ with $\left\lvert B\right\rvert\leq\frac{C\ell}{\gamma\log^{(k)}(\ell)}$ such that $\Pr[\pi|_{B}=1]\geq 1-\gamma$.
@@
There exist universal constants $\ell_{0},C$ such that for all $\ell\in\mathbb{N}$ where $\ell\geq\ell_{0}$, the following holds: Let $\pi$ be a $k$-round coin flipping protocol where $k\leq\log^{*}(\ell)-C$. Then, there exists a a set of bad players $B\subset[\ell]$ with $\left\lvert B\right\rvert\leq 0.01\ell$ and $o\in\{0,1\}$ such that $\Pr[\pi|_{B}=o]\geq 0.999$.
To prove Theorem 5.1, we will primarily utilize Theorem 4.1. We will also use a simpler result that can be established by inductively applying the KKL theorem [KKL88influence]:
@@
We will maintain two kinds of bad players: First, let $B_{R},B_{H}\subset[\ell]$ be the sets of bad players that we will use to bias round $k$. Second, we let $B_{I}\subset[\ell]$ be the set of bad players that we will use to bias rounds $1$ to $k-1$ (these will be provided to us by induction). At the end, we will take the union of all these sets, and that will be the final set of bad players that we will use to bias $\pi$. Let $h,c\in[\ell]$ be parameters that we set later (for intuition, $h$ will be set to be very small and $c$ will be set to be very large).
Given $\alpha\in(\{0,1\}^{\ell})^{(k-1)}$, we write $\pi_{\alpha}(x)$ to denote the function from $\{0,1\}^{\ell}$ to $\{0,1\}$ that corresponds to the output of the protocol when the bits that the $\ell$ players output in rounds $11,\dots,k-$ 1 are given by $\alpha$ and the bits that the players output in round $k$ are given by $x.$
@@
Since $\Pr[\pi=1]\geq\gamma$, for a non-trivial fraction of $\alpha\in(\{0,1\}^{\ell})^{k-1}$, the induced function $\pi_{\alpha}:\{0,1\}^{\ell}\to\{0,1\}$ has a non-trivial probability of outputting $1$. For the family of functions consisting of all such $\pi_{\alpha}$, we use Theorem 4.1 to find a set $B_{R}$ such that for most such $\alpha$, there exists a set $H=g(\alpha)$ with $\left\lvert H\right\rvert=h$ such that $B_{R}\cup g(\alpha)$ can bias $\pi_{\alpha}$ towards $1$. We then use induction for the first time, biasing players from the first $k-1$ rounds, so that the overall fraction of $\alpha$ for which $B_{R}\cup g(\alpha)$ can bias $\pi_{\alpha}$ towards $1$ becomes at least $3/4$.
Note that the function $g:(\{0,1\}^{\ell})^{k-1}\to 2^{[\ell]}$ can be viewed as $g(\alpha)=(g_{1}(\alpha),\dots,g_{h}(\alpha))$ where each $g_{j}$ maps ($(\{0,1\}^{\ell})^{k-1}$ to $[\ell]$ and the value $g_{j}(\alpha)$ is simply the $j$-th largest element of the set $g(\alpha)$. We will inductively find bad players that can bias each of $g_{1},\dots,g_{h}$ so that over most inputs, their image lies in a set of size $c$.
We maintain sets $C_{1},\dots,C_{h}$ so that each of $g_{1},\dots,g_{h}$ outputs an element from these respective sets with high probability. Initially all these sets are set to equal $[\ell]$. We then iteratively find some $C_{j}$ such that $\left\lvert C_{j}\right\rvert\geq c$ and cut $C_{j}$ in half. Doing this decreases by a factor of $2$ the fraction of inputs $\alpha$ whose outputs fall into the sets $C_{1},\dots,C_{h}$ ; so the fraction of inputs $\alpha$ for which $B_{R}\cup g(\alpha)$ can corrupt $\pi_{\alpha}$ is halved. To fix this, we use induction for a second time and find bad players that can corrupt the first $k-1$ rounds so that this fraction again becomes at least $3/4$.
At the end of this process, we are guaranteed that $\left\lvert C_{1}\right\rvert\leq c,\dots,\left\lvert C_{h}\right\rvert\leq c$. We then once again for a third time use induction and find bad players to bias the first $k-1$ rounds so that the fraction of $\alpha$ for which $B_{R}\cup g(\alpha)$ can corrupt $\pi_{\alpha}$ is as large as desired. We collect the union of the set of bad players outputted by $g$, i.e. the union of sets $C_{1},\dots,C_{h}$, along with $B_{R}$ and the bad players from all our inductive calls and let that be our final set of bad players.
@@
      Set =$\mathcal{F}=\{\pi_{\alpha}\}_{\operatorname*{\mathbb{E}}[\pi_{\alpha}]\geq\gamma/2}$ and apply Theorem 4.1 to $\mathcal{F}$ with parameters $h,\gamma/2$ and $\delta=1/3$. Let $B_{R}$ be the set given by Theorem 4.1, and for $\alpha\in\left(\{0,1\}^{\ell}\right)^{k-1}$ let $g:\left(\{0,1\}^{\ell}\right)^{k-1}\to\left([\ell]\right)^{h}\cup\bot$ be such that (viewing $g(\alpha)$ as a set in the obvious way) $B_{R}\cup g(\alpha)$ can $(1-\gamma/2)$-bias $\pi_{\alpha}$ towards $1$ if possible, otherwise $g(\alpha)=\bot$.
      We assert that $\Pr[g\neq\bot]\geq\gamma/6$ (this will be proven in 5.4).
@@
      Let $\pi_{temp}^{(1)}$ be the $(k-1)$-round protocol such that $\pi_{temp}^{(1)}(\alpha)=1$ iff $g(\alpha)\neq\bot$. We have that $\Pr[\pi_{temp}^{(1)}=1]\geq\gamma/6$. By induction (the first use of induction mentioned in the informal overview) we can find a set of bad players, which we denote $B_{temp}^{(1)}$, so that $\Pr[\pi_{temp}^{(1)}|_{B_{temp}^{(1)}}=1]\geq(1-\gamma/6)\geq 3/4$. Add all the bad players from $B_{temp}^{(1)}$ to $B_{I}$.
@@
    ()(*)( ∗ )   For every $\alpha$, if there exists $j\in[h]$ such that $g_{j}(\alpha)\not\in C_{j}$, then $g(\alpha)=\bot$. (Equivalently, if $g(\alpha)\neq\bot$, then $g_{j}(\alpha)\in C_{j}$ for all $j\in[h].$)
    We iteratively bias players from rounds $1$ to $k-1$ so that for all $j\in[h]$, $C_{j}$ will have $\left\lvert C_{j}\right\rvert\leq c$. At the end of each iteration, we maintain that $\Pr[g\neq\bot]\geq 3/4$ (this is equivalent to having $\Pr_{\alpha}[g_{1}(\alpha)\in C_{1},\dots,g_{h}(\alpha)\in C_{h}]\geq 3/4$). This is done as follows:
@@
        Let $\pi_{temp}^{(2)}$ be the $(k-1)$-round protocol such that $\pi_{temp}^{(2)}(\alpha)=1$ iff $g(\alpha)\neq\bot$. We have that $\Pr[\pi_{temp}^{(2)}=1]\geq 1/4$. By induction (the second use of induction mentioned in the informal overview), we can find a set of bad players, which we denote $B_{temp}^{(2)}$, so that $\Pr[\pi_{temp}^{(2)}|_{B_{temp}^{(2)}}=1]\geq 3/4$. Add all the bad players from $B_{temp}^{(2)}$ to $B_{I}$.
@@
    Let $\pi_{temp}^{(3)}$ be the $(k-1)$-round protocol such that $\pi_{temp}^{(3)}(\alpha)=1$ iff $g(\alpha)\neq\bot$. We have that $\Pr[\pi_{temp}^{(3)}=1]\geq 3/4\geq\gamma/2$. By induction (the third use of induction mentioned in the informal overview), we can find a set of bad players, which we denote $B_{temp}^{(3)}$, so that $\Pr[\pi_{temp}^{(3)}|_{B_{temp}^{(3)}}=1]\geq 1-\gamma/2$. Add all the bad players from $B_{temp}^{(3)}$ to $B_{I}$.
@@
Using the reverse Markov argument from 3.3, since $\Pr[\pi=1]\geq\gamma$, we have that $\frac{\left\lvert\mathcal{F}\right\rvert}{2^{(k-1)\ell}}=\frac{\left\lvert\alpha\in(\{0,1\}^{\ell})^{(k-1)}:\operatorname*{\mathbb{E}}[\pi_{\alpha}]\geq\gamma/2\right\rvert}{2^{(k-1)\ell}}\geq\gamma/4$. Since we set $\delta=2/3$ while using Theorem 4.1, we indeed infer that $\Pr[g\neq\bot]\geq\gamma/6$. ∎

### v2

@@ 2.7.1 Multi-output Poincaré
We introduce the necessary preliminaries in Section 3. In Section 4, we prove our main technical result on jointly biasing most functions in any family of functions. We prove our main lower bound result on the resilience of coin flipping protocols in Section 5. In Section 6, we present new multiround coin flipping protocols. In Section 7, we construct one-round random selection protocols that output uniform random bits. In Section 8, we prove our impossibility results for one-round random selection. We finally conclude with some discussion and open problems in Section 9. We provide a high level comparison of our coin flipping protocol lower bound (1) to that of [RSZ02] in Appendix A.
@@ 3.1 Notation and Terminology
For $0<\varepsilon<1/2$ and $o\in\{0,1\}$, we say that $S\subset[\ell]$ can $(1-\varepsilon)$-bias a function $f:\{0,1\}^{\ell}\to\{0,1\}$ towards $o$ if there exists an adversary $\mathcal{A}_{o}:\{0,1\}^{\ell-\left\lvert S\right\rvert}\to\{0,1\}^{\left\lvert S\right\rvert}$ such that the function $g_{o}(y):=f(\mathcal{A}_{o}(y),y)$ has $\Pr_{\mathbf{y}}[g_{o}(\mathbf{y})=o]\geq 1-\varepsilon$. We will similarly use the notion that a set $S\subseteq[\ell]$ can $(1-\varepsilon)$-bias a protocol $\pi$ towards $o$.
@@
For a function $f:\{0,1\}^{\ell}\to\{0,1\}$, we sometimes write $\Pr[f=o]$” to denote $\Pr_{\mathbf{y}\sim\{0,1\}^{\ell}}[f(\mathbf{y})=o]$. Similarly, for $\pi$ a $k$-round protocol over $\ell$ players in which each player outputs one bit per round, we sometimes write $\Pr[\pi=o]$” to denote $\Pr_{\mathbf{x}\sim{(\{0,1\}^{\ell})^{k}}}[\pi(\mathbf{x})=o]$.
@@ 3.2.1 Useful Definitions
Definition 3.1 (Shannon entropy).
For a random variable $\mathbf{X}$ over $\Omega$, we define its Shannon entropy as
$H(\mathbf{X})$ $=-\sum_{x\in\Omega}\Pr[\mathbf{X}=x]\cdot\log(\Pr[\mathbf{X}=x]).$
@@ 3.2.2 Helpful Claims
3.3 Leader Election, Collective Coin Flipping, and Random Selection Protocols
3.3.1 Formalizing Protocols in the full information model
We formalize the definition of random selection protocols in the full information model. Collective coin flipping protocols and leader election protocols are special cases of such protocols where the output domain is $\{0,1\}$ and $\ell$] respectively.
Definition 3.6 (Random Selection Protocol in the full information model).
A $k$-round random selection protocol with output domain $Y$ over $\ell$ players where each player sends $r$ random bits per round is a function
@@ 3.3.1 Formalizing Protocols in the full information model
Here is how the protocol operates in the presence of a set $B\subset[\ell]$ of bad players: In round $i$, each of the players from $\ell]\setminus B$ independently outputs a uniformly random element from $\{0,1\}^{r}$. Let their collective outputs be $\alpha_{i}\in\left(\{0,1\}^{r}\right)^{[\ell]\setminus B}$. Then, depending on $\alpha_{1},\dots,\alpha_{i}$, the players in $B$ together output an element of $\left(\{0,1\}^{r}\right)^{B}$. Hence, we model the strategy of the bad players as a sequence of functions $\sigma=(\sigma_{1},\dots,\sigma_{k})$, where
@@
In this paper we will mostly consider the case in which $r=1$ and this will be the default assumption unless stated otherwise. Note that when $k=1$, the protocol $\pi$ just becomes a function over $\{0,1\}^{\ell}$; such one-round coin flipping protocols which cannot be biased by any small set of bad players are also known as resilient functions.
@@
A leader election protocol $\pi$ is a protocol in the full information model with output domain $Y=[\ell]$, the number of players the protocol is operating on. Furthermore, we say $\pi$ is $(b,\varepsilon)$-resilient if in the presence of any set $B$ of bad players with $\left\lvert B\right\rvert\leq b$, we have that $\Pr[\pi|_{B}\in B]\leq 1-\varepsilon$.
@@ 3.3.2 A Useful Claim
Fix $\gamma\in(0,1/2]$. Suppose that for every $k$ round coin flipping protocol $\pi^{\prime}$ over $\ell$ players, there exist $b$ bad players that can $(1-\gamma)$-bias $\pi$ towards a particular outcome $o\in\{0,1\}$. Then, for every (k1)(k-1)-round leader election protocol $\pi$, there exist $b$ bad players such that that a good player is elected as a leader with probability at most $2\gamma$.
@@ 4 Biasing a Family of Functions
In this section, we show that for any family of functions mapping $\{0,1\}^{\ell}\to\{0,1\}$, there exists a “common set” $B_{R}$ of coordinates in $\ell$] such that for almost every function $f$ in the family, $B_{R}$ along with a small “heavy set” of coordinates (which may depend on $f$) can together bias $f$. Moreover, crucially, neither $B_{R}$ nor the “heavy set” is too large. This result plays an essential role in our main lower bound for $k$-round coin flipping protocols, just as its simplified version, Theorem 2.2, played a crucial role in the sketch of our two-round lower bound given in Section 2.
@@
Let $G$ be the bipartite graph with left vertex set $U$ given by $\mathcal{F}$ and right vertex set $V$ consisting of all size-$r$ subsets of $\ell$], where $r=\frac{100\ell\log(1/\delta)}{\gamma\log(h)}$. $G$ contains an edge between $f\in U$ and $B_{R}\in V$ if there exists $B_{H}=B_{H}(f,B_{R})\subset[\ell]$ with $\left\lvert B_{H}\right\rvert\leq h$ such that $B_{R}\cup B_{H}$ can $(1-\gamma)$-bias $f$ towards $o_{f}$. By Lemma 4.2, the minimum left degree of this graph is at least $(1-\delta)\binom{\ell}{r}$. This implies that the average right degree of this graph is at least $(1-\delta)\left\lvert\mathcal{F}\right\rvert$, and hence, there exists $B_{R}^{*}\in V$ with degree at least $(1-\delta)\left\lvert\mathcal{F}\right\rvert$. Thus, there must exist some fixed $B_{R}^{*}$ such that for $(1-\delta)$ fraction of $f\in\mathcal{F}$, there exists $B_{H}=B_{H}(f)$ with $|B_{H}|\leq h$ such that $B_{R}^{*}\cup B_{H}$ can $(1-\gamma)$-bias $f$ towards $o_{f}$. ∎
@@ 4.1 Proving the Main Lemma
We will use the following concentration inequality regarding submartingales from [RSZ02]:
Lemma 4.3 (Lemma 9 of [RSZ02]).
@@
We will also use the following result from [RSZ02], which follows as a slight extension of [KKL88]:
@@
Let $0<\gamma<{\frac{1}{2}},0<\theta<\frac{1}{8},\ell\in\mathbb{N}$. Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be such that $\gamma\leq\Pr[f=1]\leq 1-\gamma$. If $\operatorname{\mathbf{I}}_{i}(f)\leq\theta$ for all $i\in[\ell]$, then
$\sum_{i=1}^{\ell}\operatorname{\mathbf{I}}_{i}(f)\geq\frac{\gamma\log(1/\theta)}{20}.$
@@
$\mathbf{X}_{j}=\begin{cases}\operatorname{\mathbf{I}}_{v_{j}}(f|_{v_{1},\dots,v_{j-1}})&\text{if~}\Pr[f|_{v_{1},\dots,v_{j-1}}=1]<1-\gamma\\ 1&\textrm{otherwise}\end{cases}$
@@ 5 Biasing Coin Flipping Protocols
There exist universal constants $C=10^{7},\ell_{0}\in\mathbb{N}$ such that for all $k\in\mathbb{N},\ell\in\mathbb{N},0<\gamma<1/4$, the following holds: For any $k$-round coin flipping protocol $\pi$ over $\ell\geq\ell_{0}$ players, sending $1$ bit per round, where $\Pr[\pi=1]\geq\gamma$, there exists a set of bad players $B\subset[\ell]$ with $\left\lvert B\right\rvert\leq\frac{C\ell}{\gamma\log^{(k)}(\ell)}$ such that $\Pr[\pi|_{B}=1]\geq 1-\gamma$.
@@
There exist universal constants $\ell_{0},C$ such that for all $\ell\in\mathbb{N}$ where $\ell\geq\ell_{0}$, the following holds: Let $\pi$ be a $k$-round coin flipping protocol over $\ell$ players, sending $1$ bit per round, where $k\leq\log^{*}(\ell)-C$. Then, there exists a a set of bad players $B\subset[\ell]$ with $\left\lvert B\right\rvert\leq 0.01\ell$ and $o\in\{0,1\}$ such that $\Pr[\pi|_{B}=o]\geq 0.999$.
From Theorem 5.1, we also obtain the following lower bounds regarding leader election protocols:
Corollary 5.3.
There exist universal constants $C,\ell_{0}\in\mathbb{N}$ such that for all $k,\ell\in\mathbb{N},0<\gamma<1/4$, the following holds: For any $k$-round leader election protocol $\pi$ over $\ell\geq\ell_{0}$ players, there exists a set of bad players $B\subset[\ell]$ with $\left\lvert B\right\rvert\leq\frac{C\ell\log^{(k+1)}}{\gamma\log^{(k)}(\ell)}$ such that $\Pr[\pi|_{B}\in B]\geq 1-\gamma$.
We prove this corollary in Section 5.2.
To prove Theorem 5.1, we will primarily utilize Theorem 4.1. We will also use a simpler result that can be established by inductively applying the KKL theorem [KKL88]:
@@
We will maintain two kinds of bad players: First, let $B_{R},B_{H}\subset[\ell]$ be the sets of bad players that we will use to bias round $k$. Second, we let $B_{I}\subset[\ell]$ be the set of bad players that we will use to bias rounds $1$ to k1k-1 (these will be provided to us by induction). At the end, we will take the union of all these sets, and that will be the final set of bad players that we will use to bias $\pi$. Let $h,c\in[\ell]$ be parameters that we set later (for intuition, $h$ will be set to be very small and $c$ will be set to be very large).
Given $\alpha\in(\{0,1\}^{\ell})^{(k-1)}$, we write $\pi_{\alpha}(x)$ to denote the function from $\{0,1\}^{\ell}$ to $\{0,1\}$ that corresponds to the output of the protocol when the bits that the $\ell$ players output in rounds $1,\dots,k-1$ are given by $\alpha$ and the bits that the players output in round $k$ are given by $x.$
@@
Since $\Pr[\pi=1]\geq\gamma$, for a non-trivial fraction of $\alpha\in(\{0,1\}^{\ell})^{k-1}$, the induced function $\pi_{\alpha}:\{0,1\}^{\ell}\to\{0,1\}$ has a non-trivial probability of outputting $1$. For the family of functions consisting of all such $\pi_{\alpha}$, we use Theorem 4.1 to find a set $B_{R}$ such that for most such $\alpha$, there exists a set $H=g(\alpha)$ with $\left\lvert H\right\rvert=h$ such that $B_{R}\cup g(\alpha)$ can bias $\pi_{\alpha}$ towards $1$. We then use induction for the first time, biasing players from the first k1k-1 rounds, so that the overall fraction of $\alpha$ for which $B_{R}\cup g(\alpha)$ can bias $\pi_{\alpha}$ towards $1$ becomes at least $3/4$.
Note that the function $g:(\{0,1\}^{\ell})^{k-1}\to 2^{[\ell]}$ can be viewed as $g(\alpha)=(g_{1}(\alpha),\dots,g_{h}(\alpha))$ where each $g_{j}$ maps $(\{0,1\}^{\ell})^{k-1}$ to $\ell$] and the value $g_{j}(\alpha)$ is simply the $j$-th largest element of the set $g(\alpha)$. We will inductively find bad players that can bias each of $g_{1},\dots,g_{h}$ so that over most inputs, their image lies in a set of size $c$.
We maintain sets $C_{1},\dots,C_{h}$ so that each of $g_{1},\dots,g_{h}$ outputs an element from these respective sets with high probability. Initially all these sets are set to equal $\ell$]. We then iteratively find some $C_{j}$ such that $\left\lvert C_{j}\right\rvert\geq c$ and cut $C_{j}$ in half. Doing this decreases by a factor of $2$ the fraction of inputs $\alpha$ whose outputs fall into the sets $C_{1},\dots,C_{h}$ ; so the fraction of inputs $\alpha$ for which $B_{R}\cup g(\alpha)$ can corrupt $\pi_{\alpha}$ is halved. To fix this, we use induction for a second time and find bad players that can corrupt the first k1k-1 rounds so that this fraction again becomes at least $3/4$.
At the end of this process, we are guaranteed that $\left\lvert C_{1}\right\rvert\leq c,\dots,\left\lvert C_{h}\right\rvert\leq c$. We then once again for a third time use induction and find bad players to bias the first k1k-1 rounds so that the fraction of $\alpha$ for which $B_{R}\cup g(\alpha)$ can corrupt $\pi_{\alpha}$ is as large as desired. We collect the union of the set of bad players outputted by $g$, i.e. the union of sets $C_{1},\dots,C_{h}$, along with $B_{R}$ and the bad players from all our inductive calls and let that be our final set of bad players.
@@
      Set $\mathcal{F}=\{\pi_{\alpha}\}_{\operatorname*{\mathbb{E}}[\pi_{\alpha}]\geq\gamma/2}$ and apply Theorem 4.1 to $\mathcal{F}$ with parameters $h,\gamma/2$ and $\delta=1/3$. Let $B_{R}$ be the set given by Theorem 4.1, and for $\alpha\in\left(\{0,1\}^{\ell}\right)^{k-1}$ let $g:\left(\{0,1\}^{\ell}\right)^{k-1}\to\left([\ell]\right)^{h}\cup\bot$ be such that (viewing $g(\alpha)$ as a set in the obvious way) $B_{R}\cup g(\alpha)$ can $(1-\gamma/2)$-bias $\pi_{\alpha}$ towards $1$ if possible, otherwise $g(\alpha)=\bot$.
      We assert that $\Pr[g\neq\bot]\geq\gamma/6$ (this will be proven in 5.5).
@@
      Let $\pi_{temp}^{(1)}$ be the (k1)(k-1)-round protocol such that $\pi_{temp}^{(1)}(\alpha)=1$ iff $g(\alpha)\neq\bot$. We have that $\Pr[\pi_{temp}^{(1)}=1]\geq\gamma/6$. By induction (the first use of induction mentioned in the informal overview) we can find a set of bad players, which we denote $B_{temp}^{(1)}$, so that $\Pr[\pi_{temp}^{(1)}|_{B_{temp}^{(1)}}=1]\geq(1-\gamma/6)\geq 3/4$. Add all the bad players from $B_{temp}^{(1)}$ to $B_{I}$.
@@
    ()(*)   For every $\alpha$, if there exists $j\in[h]$ such that $g_{j}(\alpha)\not\in C_{j}$, then $g(\alpha)=\bot$. (Equivalently, if $g(\alpha)\neq\bot$, then $g_{j}(\alpha)\in C_{j}$ for all $j\in[h].$)
    We iteratively bias players from rounds $1$ to k1k-1 so that for all $j\in[h]$, $C_{j}$ will have $\left\lvert C_{j}\right\rvert\leq c$. At the end of each iteration, we maintain that $\Pr[g\neq\bot]\geq 3/4$ (this is equivalent to having $\Pr_{\alpha}[g_{1}(\alpha)\in C_{1},\dots,g_{h}(\alpha)\in C_{h}]\geq 3/4$). This is done as follows:
@@
        Let $\pi_{temp}^{(2)}$ be the (k1)(k-1)-round protocol such that $\pi_{temp}^{(2)}(\alpha)=1$ iff $g(\alpha)\neq\bot$. We have that $\Pr[\pi_{temp}^{(2)}=1]\geq 1/4$. By induction (the second use of induction mentioned in the informal overview), we can find a set of bad players, which we denote $B_{temp}^{(2)}$, so that $\Pr[\pi_{temp}^{(2)}|_{B_{temp}^{(2)}}=1]\geq 3/4$. Add all the bad players from $B_{temp}^{(2)}$ to $B_{I}$.
@@
    Let $\pi_{temp}^{(3)}$ be the (k1)(k-1)-round protocol such that $\pi_{temp}^{(3)}(\alpha)=1$ iff $g(\alpha)\neq\bot$. We have that $\Pr[\pi_{temp}^{(3)}=1]\geq 3/4\geq\gamma/2$. By induction (the third use of induction mentioned in the informal overview), we can find a set of bad players, which we denote $B_{temp}^{(3)}$, so that $\Pr[\pi_{temp}^{(3)}|_{B_{temp}^{(3)}}=1]\geq 1-\gamma/2$. Add all the bad players from $B_{temp}^{(3)}$ to $B_{I}$.
@@
Using the reverse Markov argument from 3.4, since $\Pr[\pi=1]\geq\gamma$, we have that $\frac{\left\lvert\mathcal{F}\right\rvert}{2^{(k-1)\ell}}=\frac{\left\lvert\alpha\in(\{0,1\}^{\ell})^{(k-1)}:\operatorname*{\mathbb{E}}[\pi_{\alpha}]\geq\gamma/2\right\rvert}{2^{(k-1)\ell}}\geq\gamma/4$. Since we set $\delta=1/3$ while using Theorem 4.1, we indeed infer that $\Pr[g\neq\bot]\geq\gamma/6$. ∎